Sterling silver is 92.5% pure silver. How many grams of sterling silver must be mixed to a 90% alloy to obtain a 500g of a 91% silver alloy?
Answer by jorel1380(3719)

500x.91=455
m+n=500
.9m+.925n=455
900m+925n=455000
900m+900n=450000
25n=5000
n=200
m=300
You need 300g of 90% alloy to mix with 200g of sterling silver to get your desired alloy..